Opened 12 years ago

Closed 10 years ago

Last modified 10 years ago

#29 closed flight processing (fixed)

GB2007/07, flight day 102a/2007, Inverclyde

Reported by: mggr Owned by: benj
Priority: alpha 4 high Milestone: 2007 data processing completion
Component: Processing: general Keywords:
Cc: sbg Other processors:

Description (last modified by mggr)

Data location: ~arsf/arsf_data/in_progress/2007/flight_data/uk/GB2007_07-2007_102a_Inverclyde

Data arrived from ARSF via second SATA disk transfer at the start of Sept.

Scientific details:

  • Hyperspectral monitoring of heather uplands
  • ~/arsf_data/in_progress/2007/ARSF_Applications-GB_2007/GB07-07_Application_and_science_case.pdf

Related tickets: #52

Sensors

  • Eagle (delivered 12/Mar/08)
  • Hawk (delivered 30/Apr/08)

Notes

Alasdair Mac Arthur (PI is Tim Malthus, but Alasdair is the person doing the work) has volunteered to be the 2007 Specim guinea pig.

There are several flights in this project, see: #52

Change History (24)

comment:1 Changed 12 years ago by mggr

  • Status changed from new to assigned

We're waiting on copies of the radiometric and boresight calibration (day 061) before we can begin processing. I've told him the status today and said we'll get back to him when we're able to start processing.

comment:2 Changed 12 years ago by mggr

Got radiometric now, downloading boresight calibration.

comment:3 Changed 12 years ago by mggr

  • Cc amro removed
  • Priority changed from medium to immediate

Boresight calibration is now (mostly) understood. Begun processing of this data, though will need to do another boresight prior to delivery.

Alasdair would like the data before the end of Sept to meet an evaluation license he has for some atmospheric correction software. Going to hurry this one.

comment:4 Changed 12 years ago by mggr

Processed most files (though boresight lacking). Problem with line 1 - seems like the nav file might be corrupt. Asked Bill for advice.

comment:5 Changed 12 years ago by mggr

Delivered (via FTP) some time-corrected imagery. Boresight correction remains to be done.

Apparently the site was also flown over on 10th July 2007 and 29th August 2007.

comment:6 Changed 11 years ago by mggr

  • Component changed from Support to Processing: general
  • Description modified (diff)
  • Priority changed from immediate to medium
  • Summary changed from Support: 15/June/2007, Alasdair Mac Arthur, GB2007/07 to GB2007/07, flight day 102a/2007, Inverclyde

Changing this to a normal processing ticket now we have an established standard.

Waiting for Alasdair to look at the data.

comment:7 Changed 11 years ago by mggr

  • Description modified (diff)
  • Type changed from support to flight processing

Typo fix in description.

comment:8 Changed 11 years ago by mggr

Alasdair seems to be happy with the data so far. Once we have boresight numbers confirmed, we can send them to him (or resend the data with corrections) and close this ticket.

comment:9 Changed 11 years ago by mggr

This flight has binning {4,2} (2 spatial) - may be affected by a calibration issue.

comment:10 Changed 11 years ago by anee

  • Owner changed from mggr to benj
  • Status changed from assigned to new

Processed Eagle and Hawk data located in: /users/rsg/anee/scratch_space/GB2007_07-2007_102a_Inverclyde

Delivery directory created and filled.

comment:11 Changed 11 years ago by anee

data checked by benj

synced to ~/arsf_data/in_progress/2007/flight_data/uk/GB2007_07-2007_191a_Inverclyde/delivery/20080303/GB07-07

comment:12 Changed 11 years ago by anee

  • Owner changed from benj to anee

comment:13 Changed 11 years ago by anee

synced to
/users/rsg/arsf/arsf_data/in_progress/2007/flight_data/uk/GB2007_07-2007_102a_Inverclyde/delivery/20080229/GB07-07

comment:14 Changed 11 years ago by anee

Eagle data sent to Dr. Malthus

comment:15 Changed 11 years ago by anee

  • Owner changed from anee to benj

Processed Hawk data located in: /users/rsg/anee/scratch_space/GB2007_07-2007_102a_Inverclyde
(20080414)

Delivery directory created.

comment:16 Changed 11 years ago by benj

  • Owner changed from benj to anee

Some small updates to the readme, otherwise ready to go. Rsynced to ~arsf/arsf_data/in_progress/2007/flight_data/uk/GB2007_07-2007_102a_Inverclyde.

comment:17 Changed 11 years ago by mggr

Prepared for archiving (DTG, NEODC) - moved to ~arsf/arsf_data/complete-ready_to_archive/

comment:18 Changed 11 years ago by mggr

  • Description modified (diff)

comment:19 Changed 11 years ago by anee

  • Owner changed from anee to mggr

comment:20 Changed 10 years ago by benj

  • Owner changed from mggr to mark1

Eagle and Hawk originally delivered using wrong boresight values. Re-delivering to fix this. New delivery directory created at ~benj/scratch/GB2007_07-2007_102a_Inverclyde/delivery/20081009/GB07-07/

No screenshots included with the delivery - needs to be delivered quickly and the PI has seen the previous data anyway.

comment:21 Changed 10 years ago by mark1

  • Owner changed from mark1 to benj

Delivery has been checked and looks ok. Example commands tested without dem.
Data has been rsynced over to ~arsf/arsf_data/in_progress/2007/flight_data/uk/GB2007_07-2007_102a_Inverclyde

comment:22 Changed 10 years ago by benj

  • Resolution set to fixed
  • Status changed from new to closed

Check was fine, re-delivered via FTP on 09/10/08 (but forgot to update/close ticket).

comment:23 Changed 10 years ago by mggr

Data sent to NEODC for archiving.

comment:24 Changed 10 years ago by mggr

Archived at NEODC.

Note: See TracTickets for help on using tickets.